OpenStudy (anonymous):

I dont understand this problem: Write an equation for a rational function with: Vertical asymptotes at x = -3 and x = -6 x intercepts at x = 3 and x = 4 y intercept at 3 My answer was: (3(x-3)(x-4))/((x+3)(x+6)) but the answer is: (4.5(x-3)(x-4))/((x+3)(x+6))

OpenStudy (anonymous):

vertical asmptotes at -3 and -6 means your denominator has to look like \[(x+3)(x+6)\]

OpenStudy (anonymous):

oh good you got that

OpenStudy (anonymous):

yeah, the part I dont understand is the y intercept.

OpenStudy (anonymous):

got the x intercept right too

OpenStudy (anonymous):

now you have to make sure that when x = 0, y = 3 so replace x by 0 in this \[\frac{a(x-3)(x-4)}{(x+3)(x+6)}\] set = 0 and solve for a

OpenStudy (anonymous):

i get \[\frac{a(0-3)(0-4)}{(0+3)(0+6)}=3\]

OpenStudy (anonymous):

\[a\times \frac{12}{18}=3\] \[a=3\times \frac{3}{2}\] \[a=\frac{9}{2}=4.5\]

OpenStudy (anonymous):

i meant set = 3, not set = 0 sorry

OpenStudy (anonymous):

gotcha, thank you